OpenStudy (anonymous):

I'm trying to help my brother with his math. Its a perimeter, length, width problem. I can get all the way to the equation correctly. The book even gave me the answer but I can not figure out how they got the answer. The equation is: 2(9.7+w)+2w=91.4. Find the width.

OpenStudy (anonymous):

19.4w+2w+2w=91.4 w(19.4+2+2)=91.4 w(23.4)=91.4 w=39

OpenStudy (anonymous):

after simplifying u will get 4w + 19.4 = 91.4 4w = 72 w = 18

OpenStudy (anonymous):

I thought u can't add, subtract multiply and divide unless they all have the "w" in common?

OpenStudy (anonymous):

u can since "w" is just a number

OpenStudy (anonymous):

dhanamkoilraj is right i added a w and it was their

OpenStudy (anonymous):

yes you can but before you divide you have to make sure their is nothing added or subtracted from it that's why you try to keep the variable by it self in one side

OpenStudy (anonymous):

Just a case of solving for w. dhanamkoilraj is correct.

OpenStudy (anonymous):

Hbut I do the parenthesis first....right?!?

OpenStudy (anonymous):

you foil the parenthesis, then add together the like terms of 2w+2w. then just solve for w.

OpenStudy (anonymous):

Look up the order of operations for equations. it will help!

OpenStudy (anonymous):

yea i did the parenthesis first then just added normally and equated

OpenStudy (anonymous):

Ok let me try this again...thanks so much! :)

OpenStudy (anonymous):

no probs :)

OpenStudy (anonymous):

Ok. I'm stuck as to where u came up with the second 2w at the beginning of the equation

OpenStudy (anonymous):

he distributed the 2 that was outside the parenthesis to the numbers inside the parenthesis

OpenStudy (anonymous):

she/he

OpenStudy (anonymous):

2(9.7 + w) +2w = 91.4 (2*9.7 + 2w ) + 2w = 91.4 19.4 + 2w + 2w = 91.4 u solve after this and @bere1 im a guy XD

OpenStudy (anonymous):

hehehe okay lol is cause i was sure Xp

OpenStudy (anonymous):

Oh , I thin I finally get it. :) thanks for ur patience.

OpenStudy (anonymous):

:)